# An All-Atom Generative Model for Designing Protein Complexes

### v1

@@ Abstract
Proteins typically exist in complexes, interacting with other proteins or biomolecules to perform their specific biological roles. Research on single-chain protein modeling has been extensively and deeply explored, with advancements seen in models like the series of ESM and AlphaFold. Despite these developments, the study and modeling of multi-chain proteins remain largely uncharted, though they are vital for understanding biological functions. Recognizing the importance of these interactions, we introduce APM (All-Atom Protein Generative Model), a model specifically designed for modeling multi-chain proteins. By integrating atom-level information and leveraging data on multi-chain proteins, APM is capable of precisely modeling inter-chain interactions and designing protein complexes with binding capabilities from scratch. It also performs folding and inverse-folding tasks for multi-chain proteins. Moreover, APM demonstrates versatility in downstream applications: it achieves enhanced performance through supervised fine-tuning (SFT) while also supporting zero-shot sampling in certain tasks, achieving state-of-the-art results. Code will be released at https://github.com/bytedance/apm.
@@ 1 Introduction
Figure 2: Overview of APM. APM consists of three modules: (1) A flow-matching based Seq&BB Module for generating backbone structure and sequence simultaneously; (2) a Sidechain Module for generating the all-atom structure based on the previous module’s generation; (3) A Refine Module adjusts the sequence and structure with all-atom information. The iterative denoising process enables the generation of multi-chain proteins with all-atom structure. The detailed structure of each module are presented below.
@@ 2 Related Work
The breakthrough achievements in protein structure prediction, marked by AlphaFold series (AlphaFold1-3 (Senior et al., 2020; Jumper et al., 2021; Abramson et al., 2024) and RoseTTAFold (Baek et al., 2021; Krishna et al., 2024), have revolutionized the field of protein science. With these developments, protein language models (Rives et al., 2019; Madani et al., 2020) have emerged as powerful tools. The series of ESM (Rives et al., 2019; Lin et al., 2023; Hayes et al., 2025), trained on large-scale protein sequence data, have demonstrated remarkable capabilities in protein understanding and generation. Meanwhile, certain methods considered protein design workflow in two stages: RFdiffusion (Watson et al., 2023) tackles backbone structure generation using diffusion models, while ProteinMPNN (Dauparas et al., 2022) specializes in sequence design through message-passing neural networks. FrameFlow (Yim et al., 2024, 2023b) and FoldFlow (Bose et al., ) present developments applying SE(3) flow matching approaches to protein structure generation. FoldFlow2 (Huguet et al., 2024) further demonstrates the integration of protein language models for structure generation. Besides, Chroma (Ingraham et al., 2023) introduces a unified approach to protein design through a generative model that can directly sample novel protein structures and sequences while being conditioned to target specific properties and functions The field has also seen approaches like Multiflow (Campbell et al., 2024), ProteinGenerator  (Lisanza et al., 2024), and Protpardelle (Chu et al., 2024a), which enable generation of both sequence and structure. More recently, SaProt (Su et al., 2024a, b, c) and DPLM series (Zheng et al., 2023; Wang et al., 2024a, b; Hsieh et al., 2025) have further advanced protein token modeling by incorporating structural information into the pre-training process, enabling better understanding of protein sequences.
@@
Target-specific protein design has made remarkable advances recently. In antibody design, approaches like HERN (Jin et al., 2022), DiffAb (Luo et al., 2022), MEAN (Kong et al., 2023a), and dyMEAN (Kong et al., 2023b) have demonstrated the ability to generate functional antibodies. Besides,  Wu & Li (2024),  Zhu et al. (2024), and Gao et al. (2023) introduced pre-trained protein language models as sequence priors to improve antibody design. For peptide design, methods such as PPFlow (Lin et al., 2024), PepFlow (Li et al., 2024), and PepGLAD (Kong et al., 2024) focus on designing bioactive peptides.
@@ 3.1 Representation for Protein All-Atom Structure
The all-atom protein structure is represented as a collection of amino acid types, backbone frames, and sidechain torsion angles (Jumper et al., 2021; Lehninger et al., 2005). A multi-chain protein complex $\mathcal{P}$ is composed of $K$ chains and $N=\sum_{k}k_{N}$ residues in total. For the $k$-th chain, the amino acid sequence is denoted as $\mathbf{S}_{k}=[S_{k}^{k_{1}},S_{k}^{k_{2}},\ldots,S_{k}^{k_{N}}]$, where $S_{k}^{k_{i}}\in\mathcal{A}$ and $\mathcal{A}$ is the set of 20 standard amino acids. Meanwhile, the backbone structure of this chain is characterized by rigid frames $\mathbf{T}_{k}=\left[T_{k}^{k_{1}},T_{k}^{k_{2}},\ldots,T_{k}^{k_{N}}\right]$, where each $T_{k}^{k_{i}}\in\mathrm{SE}(3)$ consists of a rotation $R_{k}^{i}\in\mathrm{SO}(3)$ and a translation vector $\mathbf{x}_{k}^{k_{i}}\in\mathbb{R}^{3}$, mapping rigid transformations from ideal peptide geometry (Engh & Huber, 2006). The sidechain torsion angles are denoted as $\mathbf{\chi}_{k}=[\chi_{k}^{k_{1}},\chi_{k}^{k_{2}},\ldots,\chi^{k_{N}}]$, where $\chi_{k}^{k_{i}}\in[0,2\pi)^{4}$ corresponds to the torsions of rotatable bonds in the sidechain of $i$-th residue. For brevity, we slightly abuse the notation such that $\mathbf{S}=\bigcup_{k}\mathbf{S}_{k}$, $\mathbf{T}=\bigcup_{k}\mathbf{T}_{k}$, $\mathbf{\chi}=\bigcup_{k}\mathbf{\chi}_{k}$, where chain indices (i.e., $k$) are omitted hereafter unless needed.
@@ 4.2 Single-Chain Protein Related Tasks
While APM is specifically designed for modeling multi-chain proteins, it also possesses the capabilities of those foundation models designed for single-chain proteins, including folding and inverse-folding. We validated the folding and inverse-folding capabilities of APM on a PDB date split used by MultiFlow. We compared it with specialized models, including ESMFold and PorteinMPNN, as well as co-design models capable of performing multiple tasks, including ESM3 and MultiFlow*(without distillation). We utilize RMSD and TMscore (Zhang & Skolnick, 2005) between predicted and ground truth structures to evaluate folding performance, and self-consistency (Trippe et al., 2022) TMscore (scTM) and amino acid recovery (AAR) to evaluate inverse-folding performance. The results are shown in Table 1.
@@
Method RMSD $\downarrow$ TM $\uparrow$ scTM $\uparrow$ AAR(%) $\uparrow$
ESMFold 2.84/1.19 0.93/0.97 - -
ProteinMPNN - - 0.94/0.97 46.58/46.76
ESM3(1.4B) 4.94/2.28 0.87/0.93 0.94/0.97 49.50/49.42
MultiFlow* 15.64/16.08 0.53/0.49 0.94/0.96 37.74/37.59
APM 4.90/2.65 0.86/0.92 0.94/0.97 50.44/50.41
@@ 4.4 Downstream Tasks
We further verify the capacity of APM on specific tasks including antibody design and peptide design, in both supervised fine-tuning (SFT) and zero-shot manner. Details about SFT refer to Sections B.4, D.3 and D.4.
@@ D.1 Sequence Sampling
D.2 Multi-Chain Protein

### v2

@@ Abstract
Proteins typically exist in complexes, interacting with other proteins or biomolecules to perform their specific biological roles. Research on single-chain protein modeling has been extensively and deeply explored, with advancements seen in models like the series of ESM and AlphaFold2. Despite these developments, the study and modeling of multi-chain proteins remain largely uncharted, though they are vital for understanding biological functions. Recognizing the importance of these interactions, we introduce APM (All-Atom Protein Generative Model), a model specifically designed for modeling multi-chain proteins. By integrating atom-level information and leveraging data on multi-chain proteins, APM is capable of precisely modeling inter-chain interactions and designing protein complexes with binding capabilities from scratch. It also performs folding and inverse-folding tasks for multi-chain proteins. Moreover, APM demonstrates versatility in downstream applications: it achieves enhanced performance through supervised fine-tuning (SFT) while also supporting zero-shot sampling in certain tasks, achieving state-of-the-art results. We released our code at https://github.com/bytedance/apm.
@@ 1 Introduction
Figure 2: Overview of APM. APM consists of three modules: (1) A flow-matching based Seq&BB Module for generating backbone structure and sequence simultaneously; (2) a Sidechain Module for generating the all-atom structure based on the previous module’s generation; (3) A Refine Module adjusts the sequence and structure with all-atom information. The iterative denoising process enables the generation of multi-chain proteins with all-atom structure. The detailed architecture of each module are presented below.
@@ 2 Related Work
The breakthrough achievements in protein structure prediction, marked by AlphaFold series (AlphaFold1-3 (Senior et al., 2020; Jumper et al., 2021; Abramson et al., 2024) and RoseTTAFold (Baek et al., 2021; Krishna et al., 2024), have revolutionized the field of protein science. With these developments, protein language models (Rives et al., 2019; Madani et al., 2020) have emerged as powerful tools. The series of ESM (Rives et al., 2019; Lin et al., 2023; Hayes et al., 2025), trained on large-scale protein sequence data, have demonstrated remarkable capabilities in protein understanding and generation. Meanwhile, certain methods considered protein design workflow in two stages: RFdiffusion (Watson et al., 2023) tackles backbone structure generation using diffusion models, while ProteinMPNN (Dauparas et al., 2022) specializes in sequence design through message-passing neural networks. FrameFlow (Yim et al., 2024, 2023b) and FoldFlow (Bose et al., 2023) present developments applying SE(3) flow matching approaches to protein structure generation. FoldFlow2 (Huguet et al., 2024) further demonstrates the integration of protein language models for structure generation. Besides, Chroma (Ingraham et al., 2023) introduces a unified approach to protein design through a generative model that can directly sample novel protein structures and sequences while being conditioned to target specific properties and functions The field has also seen approaches like Multiflow (Campbell et al., 2024), ProteinGenerator  (Lisanza et al., 2024), and Protpardelle (Chu et al., 2024a), which enable generation of both sequence and structure. More recently, SaProt (Su et al., 2024a, b, c) and DPLM series (Zheng et al., 2023; Wang et al., 2024a, b; Hsieh et al., 2025) have further advanced protein token modeling by incorporating structural information into the pre-training process, enabling better understanding of protein sequences.
@@
Target-specific protein design has made remarkable advances recently. In antibody design, approaches like HERN (Jin et al., 2022), DiffAb (Luo et al., 2022), MEAN (Kong et al., 2023a), and dyMEAN (Kong et al., 2023b) have demonstrated the ability to generate functional antibodies. Besides,  Wu & Li (2024),  Zhu et al. (2024), and Gao et al. (2023) introduced pre-trained protein language models as sequence priors to improve antibody design. For peptide design, methods such as PPFlow (Lin et al., 2024), PepFlow (Li et al., 2024), PepGLAD (Kong et al., 2024) and CpSDE (Zhou et al., 2025) focus on designing bioactive peptides.
@@ 3.1 Representation for Protein All-Atom Structure
The all-atom protein structure is represented as a collection of amino acid types, backbone frames, and sidechain torsion angles (Jumper et al., 2021; Lehninger et al., 2005). A multi-chain protein complex $\mathcal{P}$ is composed of $K$ chains and $N=\sum_{k}k_{N}$ residues in total. For the $k$-th chain, the amino acid sequence is denoted as $\mathbf{S}_{k}=[S_{k}^{k_{1}},S_{k}^{k_{2}},\ldots,S_{k}^{k_{N}}]$, where $S_{k}^{k_{i}}\in\mathcal{A}$ and $\mathcal{A}$ is the set of 20 standard amino acids. Meanwhile, the backbone structure of this chain is characterized by rigid frames $\mathbf{T}_{k}=\left[T_{k}^{k_{1}},T_{k}^{k_{2}},\ldots,T_{k}^{k_{N}}\right]$, where each $T_{k}^{k_{i}}\in\mathrm{SE}(3)$ consists of a rotation $R_{k}^{k_{i}}\in\mathrm{SO}(3)$ and a translation vector $\mathbf{x}_{k}^{k_{i}}\in\mathbb{R}^{3}$, mapping rigid transformations from ideal peptide geometry (Engh & Huber, 2006). The sidechain torsion angles are denoted as $\mathbf{\chi}_{k}=[\chi_{k}^{k_{1}},\chi_{k}^{k_{2}},\ldots,\chi^{k_{N}}]$, where $\chi_{k}^{k_{i}}\in[0,2\pi)^{4}$ corresponds to the torsions of rotatable bonds in the sidechain of $i$-th residue. For brevity, we slightly abuse the notation such that $\mathbf{S}=\bigcup_{k}\mathbf{S}_{k}$, $\mathbf{T}=\bigcup_{k}\mathbf{T}_{k}$, $\mathbf{\chi}=\bigcup_{k}\mathbf{\chi}_{k}$, where chain indices (i.e., $k$) are omitted hereafter unless needed.
@@ 4.2 Single-Chain Protein Related Tasks
While APM is specifically designed for modeling multi-chain proteins, it also possesses the capabilities of those foundation models designed for single-chain proteins, including folding and inverse-folding. We validated the folding and inverse-folding capabilities of APM on a PDB date split used by MultiFlow. We compared it with specialized models, including ESMFold and PorteinMPNN, as well as co-design models capable of performing multiple tasks, including ESM3 and MultiFlow*(without distillation). We utilize RMSD and TMscore (Zhang & Skolnick, 2005) between predicted and ground truth structures to evaluate folding performance, and self-consistency (Trippe et al., 2022) TMscore (scTM), amino acid recovery (AAR) and perplexity to evaluate inverse-folding performance. The perplexity (ppl) is provided by ProGen2-base (Madani et al., 2020; Nijkamp et al., 2023). The results are shown in Table 1.
@@
Method RMSD $\downarrow$ TM $\uparrow$ scTM $\uparrow$ AAR(%) $\uparrow$ ppl $\downarrow$
ESMFold 2.84/1.19 0.93/0.97 - - -
ProteinMPNN - - 0.94/0.97 46.58/46.76 11.44/11.48
ESM3(1.4B) 4.71/2.27 0.83/0.91 0.94/0.97 49.50/49.42 8.64/7.90
MultiFlow* 15.64/16.08 0.53/0.49 0.94/0.96 37.74/37.59 10.86/10.94
APM 4.83/2.64 0.86/0.91 0.94/0.97 50.44/50.41 8.74/8.10
@@ 4.4 Downstream Tasks
We further verify the capacity of APM on specific tasks including antibody design and peptide design, in both supervised fine-tuning (SFT) and zero-shot manner. Details about SFT refer to Sections B.4, D.4 and D.5.
@@ 5 Discussions
Impact Statement
Our work on multi-chain protein generation can be used in developing potent therapeutic macromolecues such as antibodies and accelerating the research process of drug discovery. Our method may be adapted to other scenarios of computer-aided design, such as small molecule design, material design, and chip design. It is also needed to ensure the responsible use of our method and refrain from using it for harmful purposes.
Acknowledgements
We thank anonymous reviewers for their insightful feedback. We would like to especially thank Dr. Hang Li for insightful discussions on the project and feedback on the manuscript that help shape this study. We thank Yi Zhou, Yuning Shen, Xinyou Wang, Yiming Ma, Fei Ye, and Lihao Wang for their valuable comments.
@@ B.4 Training
Protein Language Model.
We used a drop-in replacement for the ESM protein language model implementation named FAESM (Fred Zhangzhi Peng & contributors, 2024). We thank the authors for providing an efficient FlashAttention-based (Dao et al., 2022) implementation, which significantly accelerated the training speed.
@@ D.1 Sequence Sampling
D.2 Statistical Validation on Folding
We conducted folding with APM and ESM3 using 20 random seeds. For RMSD, ESM3 shows a marginally better mean (4.708±0.094 vs 4.828±0.077) with statistical significance (p $<$ 0.05). For TM-score, APM achieves better performance (0.856±0.002 vs 0.828±0.002) with statistical significance (p $<$ 0.05). The detailed results are shown below with the format of (average±std).
Table 7: Folding performance comparison between ESM3 and APM
Method RMSD  $\downarrow$ TM  $\uparrow$
ESM3 (1.4B) 4.708±0.094 0.828±0.002
APM 4.828±0.077 0.856±0.002
D.3 Multi-Chain Protein